John Smith is a single taxpayer with an AGI of 162,000. Complete Mr. Smith’s Schedule A given the following deductions: 1. Medical and dental expenses = 2,400 2. Real estate taxes from rental properties = 1,680 3. State income taxes withheld on W-2 = 6,700 4. Real estate taxes from personal residence = 3,550 5. Mortgage interest = 15,800 6. Investment interest = 130 7. Cash charitable contributions = 1,770 8. Noncash charitable contributions = 10,760 9. Unreimbursed employee expenses = 3,759 10. Tax preparation fees = 250 11. Subscriptions for business publications = 410 12. General sales tax = 2,500
Answer #1

Schedule A:

1) Medical and dental expenses=2400, AGI=162000 * 10%=1620

According to IRS you can deduct only the part of your medical and dental expense that exceeds 10% of the amount of your AGI

So here 2400-1620=780 would be allowed as deduction.

2)State and local taxes are allowed subject to a limit of 10000$. And there is a option available with the tax payer either to select income taxes or sales tax for deduction. Both cannot be applied together. So you can select anyone from the two. We here select state income taxes and not sales tax. So the deduction would be,

=6700.

Real estate taxes from personal residence=3550.

Total state and local taxes subject to a limit of 10000$ = 6700+3550=10250$. so allowed only 10000$.

Real estate taxes on rental income shall be allowed on Schedule E and not on Schedule A .

3) Mortgage interest and investment interest=15800+130=15930$.

4)Cash charitable contributions and non cash charitable contributions=1770+10760=12530$.

5) You can deduct unreimbursed employee expenses only if you are covered as an armed forces of the U.S. if you are in the army,navy,marine corps,air force,coast guard reserve,the army national guard of the U.S.or the reserve corps of the public health service. Here no information is given whether Mr. Smith is covered under the above,so we assume that he is covered and the deduction would be allowed to him=3759$

6) Tax preparation fees and subscriptions to business publications subject to 2% of AGI=162000*2%=3240. So 250 +410=660$ would be allowed as a deduction.

Total deductions=780+10000+15930+12530+3759+660=43,659$.